OpenStudy (anonymous):

Write as an algebraic expression in x: sin( arctan(3/4) )

OpenStudy (freckles):

this isn't something that varies it is just a number do you mean rewrite sin(arctan(x)) as an algebraic expression in terms of x?

OpenStudy (freckles):

then use that algebraic expression to evaluate sin(arctan(x)) for x=3/4

OpenStudy (freckles):

anyways to write sin(arctan(x)) as an algebraic expression let u=arctan(x) so tan(u)=x draw a right triangle then use that triangle to find sin(u) in terms of x

OpenStudy (anonymous):

sin(u)=3

OpenStudy (freckles):

that's impossible

OpenStudy (anonymous):

why? tan=sin/cos, so if tan=3/4, sin=3

OpenStudy (freckles):

sin is a number between -1 and 1 so sin being 3 is very impossible

OpenStudy (anonymous):

hm.

OpenStudy (freckles):

tan(u)=x is the same as tan(u)=x/1 draw a right triangle label one of the angles that isn't 90 deg u then label the opposite side x and the adjacent side 1 use Pythagorean theorem to find the hyp

OpenStudy (freckles):

then find sin(u) you will have enough info to find sin(u) since you will have opposite and hyp

OpenStudy (anonymous):

how would you find sin(u) knowing only a 90 degree angle and 1 side?

OpenStudy (freckles):

you know 2 sides

OpenStudy (freckles):

the opp is x and the adj is 1 find the hyp in terms of x

OpenStudy (freckles):

use Pythagorean theorem

OpenStudy (freckles):

big hint: opp^2+adj^2=hyp^2 this is Pythagorean theorem

OpenStudy (freckles):

you have opp and adj given to you just take the square root of both sides

OpenStudy (anonymous):

c=sqrt(x^2+1)

OpenStudy (freckles):

hyp is yes sqrt(x^2+1)

OpenStudy (freckles):

so you have opp is x and hyp is sqrt(x^2+1) recall sin(u)=opp/hyp

OpenStudy (freckles):

\[\sin(\arctan(x))=\frac{x}{\sqrt{x^2+1}}\]

OpenStudy (freckles):

you have now wrote sin(arctan(x)) as an algebraic expression

OpenStudy (freckles):

you can use that algebraic expression to evaluate sin(arctan(3/4))

OpenStudy (anonymous):

where did the x in the numerator come from?

OpenStudy (anonymous):

got it.

OpenStudy (freckles):

remember you were suppose to draw a right triangle for tan(u)=x/1 so I asked you to labeled the opposite side x and the adjacent side 1 and then I thought you used that to find the hyp

OpenStudy (anonymous):

So 3/4=x/sqrt(x+1)

OpenStudy (freckles):

not sure where you got that equation i thought you wanted to evaluate sin(arctan(3/4)) remember we got sin(arctan(x)) =x/(sqrt(x^2+1)) just replace x with 3/4 to find sin(arctan(3/4))

OpenStudy (anonymous):

oh!

OpenStudy (freckles):

\[\sin(\arctan(\frac{3}{4}))=\frac{\frac{3}{4}}{\sqrt{(\frac{3}{4})^2+1}} \\ \]

OpenStudy (anonymous):

Is it 3/5 or 12/25? My calculator says the former, my work, the latter...

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

nevermind.